OpenStudy (happpily):

(1+1/2) / (1-3/4) = How do I solve this? Studying for a placement test, please help!

OpenStudy (mathmale):

Your best bet is probably to convert both the numerator and the denom. into improper fractions. Know that process? (1+1/2) becomes 3/2. What would (1-3/4) become?

OpenStudy (mathmale):

Then, divide the 2nd fraction into the first. Remember the "invert the divisor and multiply" rule?

OpenStudy (er.mohd.amir):

(3/2)/(1/4) as 1+1/2=3/2 and 1-3/4=1/4 now 3*4/2*1=solve it =ans

OpenStudy (happpily):

6! Thank you guys so much :-) pretty darn easy lol